OpenStudy (anonymous):

You deposit $2000 in an account earning 3% interest compounded monthly. a. How much will you have in the account in 20 years? b. How much interest will you earn?

OpenStudy (solomonzelman):

Hello! lets apply the same formula here again. except that this time you will have to tell me what to plug in for each of the variables (if you can, please do, if not, ask)/ \(\Large\color{blue}{ \displaystyle {\rm A}= {\rm P}\left(1+\frac{ {\rm r} }{ {\rm n} }\right)^{{\rm n} \times {\rm t}} }\)

OpenStudy (solomonzelman):

I will tell you, that: "compounded monthly" means that it is compounded 12 times per year (so, n=12).

OpenStudy (anonymous):

A=2000(1+.3/12)12*20?

OpenStudy (anonymous):

3% = 0.03 and you need an exponent symbol, but otherwise that's right. A=2000(1+.03/12)^(12*20)

OpenStudy (anonymous):

then just subtract the original $2000 to get the interest

OpenStudy (solomonzelman):